OpenStudy (anonymous):

If the quotient of a number and 4 is added to the difference of the same number and1/2, the result is 3/4. Find the number

OpenStudy (nenadmatematika):

x/4+(x-1/2)=3/4 x+4x-2=3 5x=5 so x=1